The price index was 92 in 2014, and the inflation rate was 8.7 percent between 2013 and 2014. The price index in 2013 was

A. 100.7.
B. 83.3.
C. 84.6.
D. 100.0.
Answer #1

We know

inflation rate=change in Price index/Initial price index*100

Change in price index=Initial price index*Inflation rate/100=8.7(92)/100=8.004

New price index=8.004+92=100.004

ans is D